OpenStudy (anonymous):

Judah and 4 of his friends shared some apples. Each of the five friends had 1 1/8 of an apple. How many apples did they eat?

OpenStudy (anonymous):

5*9/8=45/8 5 and 5/8 apples

OpenStudy (anonymous):

6

OpenStudy (anonymous):

harkirat its me Dave!! u knowhangover guy!

OpenStudy (anonymous):

1 1/8 = 9/8 5*9/8 = 45/8 = 5 5/8 apples

OpenStudy (anonymous):

yes, u hv changed the picture... was it good - i mean hangover2 ??

OpenStudy (anonymous):

yeah to strong for me :D